# Quantum-Classical Hybrid Quantized Neural Network

### v2

@@ Abstract
Here in this work, we present a novel Quadratic Binary Optimization (QBO) model for quantized neural network training, enabling the use of arbitrary activation and loss functions through spline interpolation. We introduce Forward Interval Propagation (FIP), a method designed to tackle the challenges of non-linearity and the multi-layer composite structure in neural networks by discretizing activation functions into linear subintervals. This approach preserves the universal approximation properties of neural networks while allowing complex nonlinear functions to be optimized using quantum computers, thus broadening their applicability in artificial intelligence. We provide theoretical upper bounds on the approximation error and the number of Ising spins required, by deriving the sample complexity of the empirical risk minimization (ERM) problem, from an optimization perspective. A significant challenge in solving the associated Quadratic Constrained Binary Optimization (QCBO) model on a large scale is the presence of numerous constraints. When employing the penalty method to handle these constraints, tuning a large number of penalty coefficients becomes a critical hyperparameter optimization problem, increasing computational complexity and potentially affecting solution quality. To address this, we employ the Quantum Conditional Gradient Descent (QCGD) algorithm, which leverages quantum computing to directly solve the QCBO problem. We prove the convergence of QCGD under a quantum oracle with randomness and bounded variance in objective value, as well as under limited precision constraints in the coefficient matrix. Additionally, we provide an upper bound on the Time-To-Solution for the QCBO solving process. Experimental results using a coherent Ising machine (CIM) demonstrate a $94.95\%$ accuracy on the Fashion MNIST classification task, with only 1.1-bit precision. Compared to classical quantization-aware training (QAT), post-training quantization (PTQ) methods and full precision model, our approach yields higher accuracy, faster training, demonstrating its overall efficiency. In addition, evaluations against full-precision models show significant improvements in memory efficiency and inference latency, with minimal accuracy degradation. We further validate the robustness of the spline interpolation method and analyze the impact of varying precision levels on QCGD convergence, underscoring the reliability of our framework.
@@ Introduction
However, a critical limitation of CQNN emerges in its handling of activation functions. To conform to the quadratic constraints of the Ising model Hamiltonian, CQNN approximates arbitrary activation functions (e.g., ReLU, sigmoid) with polynomials. This choice raises concerns rooted in approximation theory. As established by Leshno et al. [33], multilayer feedforward networks with non-polynomial activation functions are dense in the space of continuous functions between Euclidean spaces under the compact convergence topology, a property ensuring universal approximation capability. Polynomial activation functions, conversely, limit this expressiveness, as they form a finite-degree polynomial space incapable of approximating arbitrary continuous functions [33, 34]. In addition to the limitations in expressive power, another significant drawback arises when constructing QUBO models is, polynomial activation functions require a large number of Ising spins for degree reduction. This issue further exacerbates the challenges of applying CQNN in practical scenarios. Therefore, adapting arbitrary, non-polynomial activation functions to quantum computing frameworks remains an open challenge, requiring innovative encoding strategies or hybrid quantum-classical solutions to preserve both computational efficiency and model versatility.
@@
In this study, we present a framework for training quantized neural networks using CIM. Our approach begins by formulating the quantization training problem as a QUBO problem and then leveraging quantum algorithms to overcome the shortcomings of existing methods of training quantized neural networks. Furthermore, we derive a sample complexity bound from an optimization perspective, effectively constraining the number of spins required in the quadratic binary model. This bound lays a critical foundation for scaling to large-scale scenarios by ensuring computational feasibility and efficiency. Our experimental evaluation highlights the effectiveness of this approach across multiple dimensions. Compared to standard quantization-aware training (QAT) methods, our method achieves higher accuracy with training times reduced by several orders of magnitude. When evaluated against post-training quantization (PTQ) algorithms, it delivers superior accuracy with fewer parameter bits—achieving $94.95\%$ accuracy with only $1.1$-bit precision. Furthermore, compared to full-precision models with continuous weights, our quantized networks significantly reduce memory footprint and inference latency while maintaining competitive accuracy. These results demonstrate the potential of quantum optimization to enable efficient and high-performing deep learning models for edge deployment, and paves the way for the next generation of efficient deep learning systems.
@@ Results
Here, the sum of $k$ is taken over all the neurons in the $(l-1)$-th layer.
@@
$|f(\mathbf{x})-f_{q}(\mathbf{x})|\leq\epsilon,\quad\forall\mathbf{x}\in[0,1]^{d}$ (4)
The number of weights required by $f_{q}$ has an upper bound of $\mathcal{O}(\lambda\log^{\frac{1}{\lambda-1}+1}(1/\epsilon)(1/\epsilon)^{d/n})$, while an unquantized piecewise linear neural network $f_{u}$ with the same activation function achieving the same error has an upper bound of $\mathcal{O}(\log(1/\epsilon)(1/\epsilon)^{d/n})$. Thus, the quantized weight count exceeds the unquantized one by a factor of $\mathcal{O}(\log^{5}(1/\epsilon))$, which is significantly smaller than the lower bound for unquantized network, $\Omega(\log^{-3}(1/\epsilon)(1/\epsilon)^{d/n})$.
This fact confirms that a quantized piecewise linear neural network, regardless of the specific activation, approximates target functions with efficiency comparable to an unquantized one, incurring only a low complexity overhead. Ding et al. [38] further reveal a theoretical optimal number of discrete weight values, typically lies between 1 and 4 bits across a wide range of $d$ and $\epsilon$, indicating that effective quantization of piecewise linear neural network requires only a small number of discrete weights. The efficiency supports the use of quantized piecewise linear neural networks in quantum computing contexts, where minimal discrete states could leverage quantum parallelism for training.
While Fact 1 establishes the universal approximation capability of quantized piecewise linear neural networks, it remains an interesting question to quantify the approximation error introduced when replacing a smooth activation function with a piecewise linear one, especially in the presence of spectral normalization, which is often employed to control model complexity and improve generalizability [40]. In this work, we present the following theorem to provide the error bound induced by piecewise linear approximations in a feedforward neural network.
Theorem 1
$n=O\left(\frac{1}{\sqrt{\epsilon}}\right).$ (7)
The proof of Theorem 1 is provided in the supplementary material. This result and analysis provide insight into the trade-offs between model efficiency and approximation accuracy, particularly relevant in the context of quantum-enhanced neural network training.
@@
[…]
@@
Figure 2: (a) Feedforward neural network with piecewise linear activation function. (b) Schematic representation of forward interval propagation (FIP) in a neural network, illustrating the discretization of activation functions into linear subintervals. The diagram shows multiple input neurons (left) connected through weights $(w_{0},w_{1},\ldots,w_{m})$ and a bias term $b$ to a summation node $\Sigma$, followed by an activation function. The green highlighted intervals and values of $\beta_{i}$s indicate the specific subintervals that the activation function’s output values belong to, determining the output subinterval for the next layer, effectively capturing multi-layer composite relationships during forward propagation.
We divide the interval $[\underline{M},\overline{M}]$ as $\underline{M}=M_{0}\leq M_{1}\leq M_{2}\leq\ldots\leq M_{n}=\overline{M},$ And the neural network training problem could be formulated as the following QCBO expression:
$\sum_{x}\sum_{i=1}^{n+1}{{\beta_{a^{L}_{x}}^{(i-1)}}\cdot\mathcal{L}_{x}(M_{i-1},y_{x})}$
$s.t.$ $a^{l}_{x,j}=\sum_{i=1}^{n+1}{{\beta_{z^{l}_{x,j}}^{(i-1)}}\cdot(\sigma(M_{i-1})+\sigma(M_{i}))/2},\forall x,j,l$
$\sum_{i=1}^{n}{\beta^{(i)}_{a^{L}_{x}}}M_{i-1}\leq{a^{L}_{x}}\leq\sum_{i=1}^{n}{\beta^{(i)}_{a^{L}_{x}}}M_{i},\forall x$
$\sum_{i=1}^{n}{\beta^{(i)}_{z^{l}_{x,j}}}M_{i-1}\leq\sum_{k}{w^{l}_{jk}a^{l-1}_{x,k}+b^{l}_{j}}\leq\sum_{i=1}^{n}{\beta^{(i)}_{z^{l}_{x,j}}}M_{i},\forall x,j,l$ (11)
$\sum_{i=1}^{n}{\beta^{(i)}_{a^{L}_{x}}}=1,\forall x,\sum_{i=1}^{n}{\beta^{(i)}_{z^{l}_{x,j}}}=1,\forall x,j,l$
where $x$, $y_{x}$ represents the label, $a^{L}_{x}$ denotes the output of the last layer; $a^{l}_{x,j}$ denotes the output of the $j$-th neuron of the $l$-th layer; $C(M_{i},y_{x})$ indicates the the loss when the last layer’s output is in the $i$-th subinterval and the target is $y_{x}$. For any variable $\tau$, $\beta_{\tau}=1$ if $\tau$ lies within the $i$-th sub-interval and 00 otherwise.
@@
In addition, the non-binary variables $a^{l}_{x,j}$, $w^{l}_{jk}$, $b^{l}_{j}$ are represented using binary encoding as following: $a^{l}_{x,j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{a^{l}_{x,k}},w^{l}_{jk}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{w^{l}_{jk}},b^{l}_{j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{b^{l}_{j}}.$ In supplementary, we also provide the details of the QCBO model when spline approximation with piecewise linear functions is used.
Spin Complexity and Upper Bounds.
Table 1 summarizes the required number of ising spins for various variables in our quadratic programming model, showing that the overall number of ising spins required is $O(nWDN)$, for handling the datasets with $N$ samples and the neural network with width $W$ and depth $D$.
@@
$\beta^{(i)}_{x}$ $O(nWDN)$
$w^{l}_{jk}$ $O(DW^{2})$
$b^{l}_{j}$ $O(WD)$
$a^{l}_{x,j}$ $O(WDN)$
Linearization $O(nWDN)$
Table 1: The number of required ising spins for each variable
@@
$\mathcal{L}(\theta)=\mathbb{E}_{(x,y_{x})\sim\mathcal{D}}[C_{x}(f_{\theta}(x),y_{x})],$ (12)
where $\theta=\{w^{l},b^{l}\}_{l=1}^{L}$, $f_{\theta}(x)=a^{L}_{x}$. We solve empirical risk minimization (ERM) over $N$ i.i.d. samples $\{(x_{i},y_{i})\}_{i=1}^{N}\sim\mathcal{D}$:
$\hat{\theta}_{N}=\arg\min_{\theta\in\Theta}\hat{\mathcal{L}}_{N}(\theta)=\frac{1}{N}\sum_{i=1}^{N}C_{x_{i}}(f_{\theta}(x_{i}),y_{i}).$
Note that the number of samples $N$ can be bounded while ensuring the efficiency of our learning process, this is where the concept of VC dimension becomes crucial.
@@
The VC dimension of a neural network with piecewise polynomial activation functions is bounded by $O(W^{3}D^{2})$. Specifically, when the activation functions are piecewise constant, the VC dimension is no more than $O(W^{2}D(\log W+\log D))$.
Within the Probably Approximately Correct (PAC) learning framework [44], the focus is on controlling the generalization error [44], a measure of how accurately an algorithm is able to predict outcomes for previously unseen data. This framework ensures that, with high probability $1-\alpha$, the generalization error is confined within a precision $\varepsilon$. Consequently, the required number of samples can be bounded as:
$N=O\Big{(}\frac{\text{VCdim}}{\varepsilon}\log\frac{1}{\alpha}\Big{)}=O\Big{(}\frac{W^{2}D(\log W+\log D)}{\varepsilon}\log\frac{1}{\alpha}\Big{)}.$ (13)
@@
For a quantized neural network with $B$-bit parameters, let $\varepsilon,\alpha\in(0,1)$, and assume the loss function $C_{x}(\cdot)$ is bounded in $[0,C_{\max}]$. Then, with probability at least $1-\alpha$,
@@
provided $N=O\left(\frac{C_{\max}^{2}}{\varepsilon^{2}}\log\frac{|\Theta|}{\alpha}\right)$. Substituting $|\Theta|$:
$N=O\left(\frac{C_{\max}^{2}}{\varepsilon^{2}}\sum_{l=1}^{L}d_{l}(d_{l-1}+1)B\log 2+\log\frac{1}{\alpha}\right),$ (15)
@@
Proof: Define the function class $\mathcal{G}=\{g_{\theta}(x,y)=C_{x}(f_{\theta}(x),y):\theta\in\Theta\}$. Given the finite $\Theta$, apply Hoeffding’s inequality for $g_{\theta}\in[0,C_{\max}]$:
$P\left(\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2\exp\left(-\frac{2N\varepsilon^{2}}{C_{\max}^{2}}\right).$ (16)
@@
$P\left(\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{C_{\max}^{2}}\right).$ (17)
Set the probability to $\alpha$, i.e., $2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{C_{\max}^{2}}\right)\leq\alpha$, we have $N\geq\frac{C_{\max}^{2}}{2\varepsilon^{2}}\log\frac{2|\Theta|}{\alpha}$. Since $\hat{\theta}_{N}$ minimizes ^$\hat{\mathcal{L}}_{N}$:

### v4

@@ Abstract
In this work, we introduce a novel Quadratic Binary Optimization (QBO) framework for training a quantized neural network. The framework enables the use of arbitrary activation and loss functions through spline interpolation, while Forward Interval Propagation addresses the nonlinearities and the multi-layered, composite structure of neural networks via discretizing activation functions into linear subintervals. This preserves the universal approximation properties of neural networks while allowing complex nonlinear functions accessible to quantum solvers, broadening their applicability in artificial intelligence. Theoretically, we derive an upper bound on the approximation error and the number of Ising spins required by deriving the sample complexity of the empirical risk minimization problem from an optimization perspective. A key challenge in solving the associated large-scale Quadratic Constrained Binary Optimization (QCBO) model is the presence of numerous constraints. When employing the penalty method to handle these constraints, tuning a large number of penalty coefficients becomes a critical hyperparameter optimization problem, increasing computational complexity and potentially affecting solution quality. To overcome this, we adopt the Quantum Conditional Gradient Descent (QCGD) algorithm, which solves QCBO directly on quantum hardware. We establish the convergence of QCGD under a quantum oracle subject to randomness, bounded variance, and limited coefficient precision, and further provide an upper bound on the Time-To-Solution. To enhance scalability, we further incorporate a decomposed copositive optimization scheme that replaces the monolithic lifted model with sample-wise subproblems. This decomposition substantially reduces the quantum resource requirements and enables efficient low-bit neural network training. We further propose the usage of QCGD and Quantum Progressive Hedging (QPH) algorithm to efficiently solve the decomposed problem. Experimental results using a coherent Ising machine achieve $94.95\%$ accuracy on the Fashion MNIST classification task with only 1.1-bit precision. Compared to classical quantization-aware training, post-training quantization methods, and a full precision model, our approach delivers higher accuracy, faster training, demonstrating its overall efficiency. Additionally, evaluations show substantial improvements in memory footprint and inference latency with negligible accuracy loss. We further validate the robustness of the spline interpolation method and analyze the effect of precision levels on QCGD convergence, confirming the reliability of the proposed framework.
@@ Introduction
However, a critical limitation of CQNN emerges in its handling of activation functions. To conform to the quadratic constraints of the Ising model Hamiltonian, CQNN approximates arbitrary activation functions (e.g., ReLU, sigmoid) with polynomials. This choice raises concerns rooted in approximation theory. As established by Leshno et al. [33], multilayer feedforward networks with non-polynomial activation functions are dense in the space of continuous functions between Euclidean spaces under the compact convergence topology, a property ensuring universal approximation capability. Polynomial activation functions, conversely, limit this expressiveness, as they form a finite-degree polynomial space incapable of approximating arbitrary continuous functions [33, 34]. In addition to the limitations in expressive power, another significant drawback that arises when constructing QUBO models is, polynomial activation functions require a large number of Ising spins for degree reduction. This issue further exacerbates the challenges of applying CQNN in practical scenarios. Therefore, adapting arbitrary, non-polynomial activation functions to quantum computing frameworks remains an open challenge, requiring innovative encoding strategies or hybrid quantum-classical solutions to preserve both computational efficiency and model versatility.
@@
In this study, we present a framework for training quantized neural networks using CIM.
Our main contributions are summarized as follows.
  • Unified quantum-compatible and convex formulation for neural networks training. We develop a general spline-based discretization scheme combined with Forward Interval Propagation (FIP), which transforms arbitrary activation and loss functions into a QCBO model. By leveraging copositive programming, we show that this QCBO formulation admits an exact convex lifting, allowing (quantized and full precision) neural networks of arbitrary depth, width, and piecewise polynomial activations to be represented as a single convex optimization problem over the completely positive cone. The resulting CP formulation preserves all nonlinear and combinatorial interactions exactly, while maintaining a convex feasible region in the lifted space. Moreover, we derive a sample complexity bound from an optimization perspective, effectively constraining the number of spins required in the quadratic binary model. This bound lays a critical foundation for scaling to large-scale scenarios by ensuring computational feasibility and efficiency.
  • A scalable quantum classical optimization framework with provable robustness. We introduce the noise-robust version of Quantum Conditional Gradient Descent (QCGD) algorithm tailored to the copositive formulation and prove its convergence under realistic quantum oracles with randomness and limited coefficient precision. To further address the scalability challenge of the monolithic lifted model, we adopt a Decomposed Lower-Bound Optimization (DLBO) strategy that replaces the global cone with sample-wise copositive blocks [38, 39]. This decomposition enables efficient bit-level optimization per sample and plays a role analogous to stochastic gradient descent in classical deep learning. It substantially reduces the dimension of each quantum subproblem while preserving the tightness of the relaxation. We further show how both the QCGD algorithm and newly introduced Quantum Progressive Hedging (QPH) algorithm can be adopted to solve the decomposed problem.
  • Training and empirical validation. Our experimental evaluation highlights the effectiveness of this approach across multiple dimensions. Compared to standard quantization-aware training (QAT) methods, our method achieves higher accuracy with training times reduced by several orders of magnitude. When evaluated against post-training quantization (PTQ) algorithms, it delivers superior accuracy with fewer parameter bits, achieving $94.95\%$ accuracy with only $1.1$-bit precision. Furthermore, compared to full-precision models with continuous weights, our quantized networks significantly reduce memory footprint and inference latency while maintaining competitive accuracy. To further validate the scalability and effectiveness of the proposed DLBO-based training framework, we extend our evaluation beyond binary classification and conduct experiments on multiclass classification tasks using the Fashion-MNIST, Wine, and Digits datasets. Across all datasets and task settings, our method consistently delivers high accuracy.
These results demonstrate the potential of quantum optimization to enable efficient and high-performing deep learning models for edge deployment and pave the way for the next generation of efficient deep learning systems.
@@ Results
Here, the sum of $k$ is taken over all the neurons in the (l1)(l-1)-th layer.
@@
$|f(\mathbf{x})-f_{q}(\mathbf{x})|\leq\epsilon,\quad\forall\mathbf{x}\in[0,1]^{d}.$ (4)
The number of weights required by $f_{q}$ has an upper bound of $\mathcal{O}(\lambda\log^{\frac{1}{\lambda-1}+1}(1/\epsilon)(1/\epsilon)^{d/n})$, while an unquantized piecewise linear neural network $f_{u}$ with the same activation function achieving the same error has an upper bound of $\mathcal{O}(\log(1/\epsilon)(1/\epsilon)^{d/n})$. Thus, the quantized weight count exceeds the unquantized one by a factor of $\mathcal{O}(\log^{5}(1/\epsilon))$, which is significantly smaller than the lower bound for the unquantized network, $\Omega(\log^{-3}(1/\epsilon)(1/\epsilon)^{d/n})$.
This fact confirms that a quantized piecewise linear neural network, regardless of the specific activation, approximates target functions with efficiency comparable to an unquantized one, incurring only a low complexity overhead. Ding et al. [40] further reveal a theoretical optimal number of discrete weight values, which typically lies between 1 and 4 bits across a wide range of $d$ and $\epsilon$, indicating that effective quantization of piecewise linear neural networks requires only a small number of discrete weights. The efficiency supports the use of quantized piecewise linear neural networks in quantum computing contexts, where minimal discrete states could leverage quantum parallelism for training.
While Fact 1 establishes the universal approximation capability of quantized piecewise linear neural networks, it remains an interesting question to quantify the approximation error introduced when replacing a smooth activation function with a piecewise linear one. For completeness, we provide the error bound induced by piecewise linear approximations in a feedforward neural network in Section 2 of the supplementary material.
@@
$a^{l}_{x}$ $=\sigma(z^{l}_{x}),\quad\forall x,\;1\leq l\leq L,$ (6)
$z^{l}_{x}$ $=\mathbf{W}_{l}\cdot a^{l-1}_{x}+\mathbf{b}_{l},\quad\forall x,\;1\leq l\leq L,$ (7)
@@
Figure 2: (a) Feedforward neural network with piecewise linear activation function. (b) Schematic representation of forward interval propagation (FIP) in a neural network, illustrating the discretization of activation functions into linear subintervals. The diagram shows multiple input neurons (left) connected through weights $(w_{0},w_{1},\ldots,w_{m})$ and a bias term $b$ to a summation node $\Sigma$, followed by an activation function. The green-highlighted intervals and values of $\beta_{i}$s indicate the specific subintervals that the activation function’s output values belong to, determining the output subinterval for the next layer, effectively capturing multi-layer composite relationships during forward propagation.
We divide the interval $\underline{M},\overline{M}]$ as $\underline{M}=M_{0}\leq M_{1}\leq M_{2}\leq\ldots\leq M_{n}=\overline{M},$ and the neural network training problem could be formulated as the following QCBO expression:
$\min$ $\sum_{x}\sum_{i=1}^{n+1}\beta_{a^{L}_{x}}^{(i-1)}\cdot\mathcal{L}(M_{i-1},y_{x})$ (8)
s.t. $a^{l}_{x,j}=\sum_{i=1}^{n+1}\beta_{z^{l}_{x,j}}^{(i-1)}\cdot\frac{\sigma(M_{i-1})+\sigma(M_{i})}{2},\quad\forall x,j,l,$ (9)
$\sum_{i=1}^{n}\beta^{(i)}_{a^{L}_{x}}M_{i-1}\leq a^{L}_{x}\leq\sum_{i=1}^{n}\beta^{(i)}_{a^{L}_{x}}M_{i},\quad\forall x,$ (10)
$\sum_{i=1}^{n}\beta^{(i)}_{z^{l}_{x,j}}M_{i-1}\leq\sum_{k}(w^{l}_{jk}a^{l-1}_{x,k}+b^{l}_{j})\leq\sum_{i=1}^{n}\beta^{(i)}_{z^{l}_{x,j}}M_{i},\quad\forall x,j,l,$ (11)
$\sum_{i=1}^{n}\beta^{(i)}_{a^{L}_{x}}=1,\quad\forall x,$ (12)
$\sum_{i=1}^{n}\beta^{(i)}_{z^{l}_{x,j}}=1,\quad\forall x,j,l,$ (13)
where $x$, $y_{x}$ represents the label, $a^{L}_{x}$ denotes the output of the last layer; $z^{l}_{x,j}$ and $a^{l}_{x,j}$ denotes the input and output of the $j$-th neuron of the $l$-th layer; $\mathcal{L}(M_{i},y_{x})$ indicates the loss when the last layer’s output is in the $i$-th subinterval and the target is $y_{x}$. For any variable $\tau$, $\beta^{(i)}_{\tau}=1$ if $\tau$ lies within the $i$-th sub-interval and 0 otherwise.
@@
In addition, the non-binary variables $a^{l}_{x,j}$, $w^{l}_{jk}$, $b^{l}_{j}$ are represented using binary encoding as following: $a^{l}_{x,j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{a^{l}_{x,k}},w^{l}_{jk}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{w^{l}_{jk}},b^{l}_{j}=\sum_{\ell}2^{\ell}\cdot\delta^{(\ell)}_{b^{l}_{j}}$. In Section 3 of supplementary, we provide further details of spline approximation.
From Multi-level Quantization to Binary [43].
For layer $\ell\in[L]$, let the trainable weight tensor be vectorized as $\mathbf{w}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ and each coordinate $p\in[d_{\ell}]$ be constrained to a finite codebook
$w^{(\ell)}_{p}\ \in\ \mathcal{S}^{(\ell)}_{p}\ \subset\ \mathbb{R},\qquad\mathcal{S}^{(\ell)}_{p}=\{\alpha^{(\ell)}_{p}\,c^{(\ell)}_{p,1}+\beta^{(\ell)}_{p},\ \dots,\ \alpha^{(\ell)}_{p}\,c^{(\ell)}_{p,K_{p}}+\beta^{(\ell)}_{p}\}$, (14)
where $\alpha^{(\ell)}_{p},\beta^{(\ell)}_{p}\in\mathbb{R}$ allow per-layer or per-channel scaling/shift and $\{c^{(\ell)}_{p,k}\}_{k=1}^{K_{p}}$ are fixed codes (not necessarily equally spaced or symmetric). Biases are treated identically.
One-hot encoding of weights. Introduce for each $(\ell,p)$ the binary indicator block
$\mathbf{y}^{(\ell)}_{p}=(y^{(\ell)}_{p,1},\dots,y^{(\ell)}_{p,K_{p}})^{\top}\in\{0,1\}^{K_{p}},\qquad\sum_{k=1}^{K_{p}}y^{(\ell)}_{p,k}=1,$ (15)
and define the affine map
$w^{(\ell)}_{p}\ =\ \sum_{k=1}^{K_{p}}\Big(\alpha^{(\ell)}_{p}\,c^{(\ell)}_{p,k}+\beta^{(\ell)}_{p}\Big)\,y^{(\ell)}_{p,k}$. (16)
Table 1 summarizes the required number of Ising spins for various variables in our quadratic programming model, showing that the overall number of Ising spins required is $O(nWLN)$, for handling the datasets with $N$ samples and the neural network with width $W$ and depth $L$.
@@
$\beta^{(i)}_{x}$ $O(nWLN)$
$w^{l}_{jk}$ $O(LW^{2})$
$b^{l}_{j}$ $O(WL)$
$a^{l}_{x,j}$ $O(WLN)$
Linearization $O(nWLN)$
Table 1: The number of required Ising spins for each variable.
@@
$\mathcal{L}(\theta)=\mathbb{E}_{(x,y_{x})\sim\mathcal{P}}[C_{x}(f_{\theta}(x),y_{x})],$ (17)
where $\theta=\{w^{l},b^{l}\}_{l=1}^{L}$, $f_{\theta}(x)=a^{L}_{x}$. Then, we solve the empirical risk minimization (ERM) over $N$ i.i.d. samples $\{(x_{i},y_{i})\}_{i=1}^{N}\sim\mathcal{P}$:
$\hat{\theta}_{N}=\arg\min_{\theta\in\Theta}\hat{\mathcal{L}}_{N}(\theta)=\frac{1}{N}\sum_{i=1}^{N}C_{x_{i}}(f_{\theta}(x_{i}),y_{i}).$ (18)
Note that the number of samples $N$ can be bounded while ensuring the efficiency of our learning process; this is where the concept of VC dimension becomes crucial.
@@
The VC dimension of a neural network with piecewise polynomial activation functions is bounded by $O(W^{3}L^{2})$. Specifically, when the activation functions are piecewise constant, the VC dimension is no more than $O(W^{2}L(\log W+\log L))$.
Within the Probably Approximately Correct (PAC) learning framework [46], we focus on controlling the generalization error, a measure of how accurately an algorithm is able to predict outcomes for previously unseen data. This framework ensures that, with high probability $1-\alpha$, the generalization error is confined within a precision $\varepsilon$. Consequently, the required number of samples can be bounded as:
$N=O\Big(\frac{\text{VCdim}}{\varepsilon}\log\frac{1}{\alpha}\Big)=O\Big(\frac{W^{2}L(\log W+\log L)}{\varepsilon}\log\frac{1}{\alpha}\Big).$ (19)
@@
For a quantized neural network with $B$-bit parameters, let $\varepsilon,\alpha\in(0,1)$, and assume the loss function $\mathcal{L}(\cdot)$ is bounded in $0,\mathcal{L}_{\max}]$. Then, with probability at least $1-\alpha$,
@@
provided $N=O\left(\frac{\mathcal{L}_{\max}^{2}}{\varepsilon^{2}}\log\frac{|\Theta|}{\alpha}\right)$. Substituting $|\Theta|$:
$N=O\left(\frac{\mathcal{L}_{\max}^{2}}{\varepsilon^{2}}\sum_{l=1}^{L}d_{l}(d_{l-1}+1)B+\log\frac{1}{\alpha}\right),$ (21)
@@
Proof: Define the function class $\mathcal{G}=\{g_{\theta}(x,y)=\mathcal{L}(f_{\theta}(x),y):\theta\in\Theta\}$. Given the finite $\Theta$, apply Hoeffding’s inequality for $g_{\theta}\in[0,\mathcal{L}_{\max}]$:
$P\left(\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2\exp\left(-\frac{2N\varepsilon^{2}}{\mathcal{L}_{\max}^{2}}\right).$ (22)
@@
$P\left(\sup_{\theta\in\Theta}\left|\frac{1}{N}\sum_{i=1}^{N}g_{\theta}(x_{i},y_{i})-\mathbb{E}_{(x,y_{i})}[g_{\theta}(x_{i},y_{i})]\right|>\varepsilon/2\right)\leq 2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{\mathcal{L}_{\max}^{2}}\right).$ (23)
Set the probability to $\alpha$, i.e., $2|\Theta|\exp\left(-\frac{2N\varepsilon^{2}}{\mathcal{L}_{\max}^{2}}\right)\leq\alpha$, we have $N\geq\frac{\mathcal{L}_{\max}^{2}}{2\varepsilon^{2}}\log\frac{2|\Theta|}{\alpha}$. Since $\hat{\theta}_{N}$ minimizes $\hat{\mathcal{L}}_{N}$: